OpenStudy (anonymous):

find the product (2/9) x (9/5)

OpenStudy (turingtest):

to multiply fractions, just multiply the numbers straight across

OpenStudy (anonymous):

that's 18/45

OpenStudy (turingtest):

correct :)

OpenStudy (turingtest):

you can simplify it from there

OpenStudy (turingtest):

do you know how to do that?

OpenStudy (anonymous):

yeah , but my answer choices are A. 11/45 B. 18/40 C. 10/81 D. 2/5

OpenStudy (turingtest):

if you know how to simplify 18/45, then you see that your answer is on that list

OpenStudy (anonymous):

okay hold on

OpenStudy (turingtest):

to simplify, find the largest number that evenly divides both the numerator and denominator

OpenStudy (anonymous):

2/5

OpenStudy (turingtest):

yep :)

OpenStudy (anonymous):

thanks man

OpenStudy (turingtest):

much easier way is to notice that you can rewrite this as (2/9) x (9/5) = (9/9)x(2/5)=2/5 saves a lot of time ;) welcome